# Discrete-Time Open Quantum Walks for Vertex Ranking in Graphs

### v2

@@ 4 Kraus operators to restart the walking in discrete-time open quantum walk
In practice, instead of calculating $t\rightarrow\infty$, we compute $q\Pi^{(t)}$ for a finite number of iterations on $t$ to achieve the limiting probability distribution. To determine the number of required iterations we fix a stopping criterion. Here, we need the distance between two qPageRank vectors. The Euclidean distance between $q\Pi^{(T-1)}$ and $q\Pi^{(T)}$ in two consiquitive steps $t=(T-1)$ and $t=T$ is determined by [42]
@@
Here, we discuss a significant benefit of utilizing the discrete-time open quantum walk for determining qPageRank. We have already mentioned that there are multiple proposals of qPageRank depending on various quantum walks; for instance, the qPageRank based on discrete-time quantum walks [16], qPageRank based on continuous-time open quantum walk [17, 18], qPageRank based on Szegedy’s quantum walk [19, 20, 21], and [22], etc. The quantum walk dynamics of the discrete-time quantum walk, the continuous-time open quantum walk, and Szegedy’s quantum walk are led by unitary matrices. At every step of the quantum walk, they calculate an instantaneous qPageRank $q\Pi^{(t)}$. The unitary matrix governing the quantum walk prevents the sequence of instantaneous qPageRank from convergent. To assure convergence, we consider the time average of instantaneous qPageRank, which is $\{\frac{1}{T}\sum_{t=0}^{T}q\Pi^{(t)}\}_{T=0}^{\infty}$ [11]. In these methods, $\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=0}^{T}q\Pi^{(t)}$ is considered as qPageRank. In our case, we involve the Markovian dynamics, which assures the convergence of $q\Pi^{(t)}$ to a limit for $t\rightarrow\infty$. Therefore, we do not need the sequence of time averages. It reduces the calculations and makes the process faster.
@@
In this section, we present our observations from the numerical experiments with Procedure 1 on a number of graphs and networks. We also compare the qPageRank with the PageRank for both directed and undirected graphs. Our construction of qPageRank is applicable to the directed graphs. Recall that we can represent an undirected graph as a directed graph by considering two opposite directions on every edge. We plot bar diagrams to represent the importance of different vertices based on PageRank and qPageRank. In all the bar diagrams, the blue bars represent classical PageRank and the orange bars represent qPageRank. Our computations utilize the Python library NetworkX [40].
The PageRank and qPageRank may behave differently depending on the graphs. To compare them we try to realize how the ranks of the vertices generated by these two methods are correlated. We rank the vertices of the graph with the integers $\{1,2,\dots n\}$ based on PageRank and qPageRank values separately. We break the ties arbitrarily if occur. Then we calculate Kendall’s rank correlation coefficient [46], if the ranking on the vertices mismatch.
@@ 6.1 Comparison between PageRank and qPageRank for a few simple graphs
(a) The Path graph with 10 vertices.
@@
Figure 1: Comparison between the PageRank and qPageRank of vertices on path graph with 10 vertices. (Color online)
@@ 6.2 Comparison between PageRank and qPageRank for a few directed graphs
Now, we consider the directed graphs, where most edges have only one direction. There are a few graphs for which PageRank and qPageRank acts similarly, such as the directed trinary tree, GN, GNC, and GNR networks [50]. For many others PageRank and qPageRank behaves differently, for instance, the scale-free network, Erdős-Rényi random graph, random_k_out graph, etc. We initiate our discussion with the graphs for which PaeRank and qPageRank match.
We first consider a directed trinary tree for our computation. Here, edges are directed outwards from a central vertex 00. The vertices are distributed into four layers surrounding the central vertex. The vertices in the same layer have equal PageRank and qPageRank, which we mention in Table 6.
@@
We plot the values of PageRank and qPageRank of the directed tree using bar diagrams. The graph and its corresponding bar diagrams are kept in Figure 3. The PageRank and qPageRank indicate the same ranking pattern on the directed trinary tree. As a result, the rank correlation is $1$. Also, the difference in PageRank of the vertices in the third and fourth layers is $0.000039$. But, the difference in qPageRank of the vertices in the third and fourth layers is .0091.0091.0091.0091. Therefore, the qPageRank is more sensitive for determining the vertices in the outer layer. Also, compare the bar diagrams in Figures 2(b) and 3(b), which are depicted for the undirected and directed trinary trees with an equal number of vertices, respectively. The qPageRank is more efficient than the classical PageRank for finding the vertices in the outermost layer in the case of a directed trinary tree.
@@
The Erdős-Rényi random graph [erdHos1960evolution] is a family of well-known random graphs. The edges are selected randomly to construct a random graph. We consider a specific instance of a directed Erdős-Rényi random graph with $80$ vertices and $283$ edges, which we plot in figure 8(a). In this case, the PageRank and qPageRank behave differently. The top five vertices with the highest PageRank are $5,79,78,45$, and $3$. Also, the top five vertices with the highest qPageRank are $26,67,77,68$, and $17$. There is no common vertex. The difference between PageRank and qPageRank for the Erdős-Rényi random graph is also talked about in other qPageRank proposals [19, 22].
@@ 6.3 Convergence test of qPageRank
For our numerical experiment to determine convergence of qPageRank procedure, we consider a random sample of the vertices from a graph, which includes the vertex with the maximum qPageRank. For different time instance $t=0,1,2,\dots T$, we calculate $q\Pi^{(t)}$ and plot corresponding values for distinct vertices in our sample. Note that, we consider accuracy of $q\Pi$ upto three decimal places. Therefore, $q\Pi^{(T)}$ represents the accurate value of $q\Pi$ upto three decimal placed. For better accuracy the number of iterations may be increased. For different graphs we collect the plots in Figure 10.
@@ 6.4 Dependency of qPageRank on $\alpha$
There are experimental evidences showing that the PageRank exhibits a dependency on the damping parameter $\alpha$. The usual choice of $\alpha=0.85$ is expected to produce a satisfactory web surfing model, as well as a good convergence rate in computing the PageRank vector. The stability of the qPageRank is also a crucial factor to consider, as it randomly sets the damping parameter to a certain value. In fact, there is no a priori argument available to set the value of $\alpha$. The initial value of $0.85$ in the conventional PageRank protocol was selected to resemble the actions of a random walker. Given the arbitrary selection of the parameter $\alpha$, it is highly preferred that the algorithm’s output remain stable. Hence, it is interesting to investigate how the qPageRank is inspired by the choice of $\alpha$. For the other proposals, the dependency of qPageRank on $\alpha$ is investigated in [16, 19, 20, 22, 18]
@@
We calculate the trace distance $D(\rho_{average},\rho^{(T)})$ and fidelity $F(\rho_{average},\rho^{(T)})$ between the average state $\rho_{average}$ and the final state $\rho^{(T)}$ for different values of $\alpha$. In figure 11 and 12, we plot $D(\rho_{average},\rho^{(T)})$ and $F(\rho_{average},\rho^{(T)})$ with respect to $\alpha$ for different undirected and directed graphs, respectively, under our consideration. Below we summaries our observations from these figures.
Consider the undirected graphs first. We observe that if the trace distance $D(\rho_{average},\rho^{(T)})$ has a local maxima for some $\alpha=\alpha_{0}\in(0,1)$, then the fidelity $F(\rho_{average},\rho^{(T)})$ has a local minima at $\alpha_{0}$. For instance, consider the graph for $D(\rho_{average},\rho^{(T)})$ and $F(\rho_{average},\rho^{(T)})$, which are depicted in sub-figures 11(a) and 11(b), respectively, for the balanced tree $(3,4)$. In sub-figure 11(a) we observe a local maxima at $\alpha=0.85$. Also, in sub-figure 11(b) we obtain a local minima at $\alpha=0.85$. Similarly, these local maxima and minima can be observed at $\alpha=0.75$ in the case of Karate club network, which are depicted in figures 11(c) and 11(d). The Barabási and Albert network has local maxima and minima for $\alpha=0.95$, which is depicted in sub-figures 11(e) and 11(f). In the case of the Watts-Strogatz network there are no such local maxima or minima. The corresponding curves are drawn in sub-figure 11(g) and 11(h). Therefore, we may assume that $\alpha=0.85$ is a wise choice.

### v3

@@ 4 Kraus operators to restart the walking in discrete-time open quantum walk
In practice, instead of calculating $t\rightarrow\infty$, we compute $q\Pi^{(t)}$ for a finite number of iterations on $t$ to achieve the limiting probability distribution. To determine the number of required iterations we fix a stopping criterion. Here, we need the distance between two qPageRank vectors. The Euclidean distance between $q\Pi^{(T-1)}$ and $q\Pi^{(T)}$ in two consecutive steps $t=(T-1)$ and $t=T$ is determined by [45]
@@
Here, we discuss a significant benefit of utilizing the discrete-time open quantum walk for determining qPageRank. We have already mentioned that there are multiple proposals of qPageRank depending on various quantum walks; for instance, the qPageRank based on discrete-time quantum walks [21], qPageRank based on continuous-time open quantum walk [22, 17], qPageRank based on Szegedy’s quantum walk [23, 24, 25], and [26], etc. The quantum walk dynamics of the discrete-time quantum walk, the continuous-time open quantum walk, and Szegedy’s quantum walk are led by unitary matrices. At every step of the quantum walk, they calculate an instantaneous qPageRank $q\Pi^{(t)}$. The unitary matrix governing the quantum walk prevents the convergence of the instantaneous qPageRank sequence. To assure convergence, we consider the time average of instantaneous qPageRank, which is $\{\frac{1}{T}\sum_{t=0}^{T}q\Pi^{(t)}\}_{T=0}^{\infty}$ [11]. In these methods, $\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=0}^{T}q\Pi^{(t)}$ is considered as qPageRank. In our case, we involve the Markovian dynamics, which assures the convergence of $q\Pi^{(t)}$ to a limit for $t\rightarrow\infty$. Therefore, we do not need the sequence of time averages. It reduces the calculations and makes the process faster.
@@
In this section, we present our observations from the numerical experiments with Procedure 1 on a number of graphs and networks. We also compare the qPageRank with the PageRank for both directed and undirected graphs. Our construction of qPageRank is applicable to directed graphs. Recall that we can represent an undirected graph as a directed graph by considering two opposite directions on every edge. We plot bar diagrams to represent the importance of different vertices based on PageRank and qPageRank. In all the bar diagrams, the blue bars represent classical PageRank and the orange bars represent qPageRank. Our computations utilize the Python library NetworkX [41].
The PageRank and qPageRank may behave differently depending on the graphs. To compare them we try to realize how the ranks of the vertices generated by these two methods are correlated. We rank the vertices of the graph with the integers $\{1,2,\dots n\}$ based on PageRank and qPageRank values separately. We break the ties arbitrarily. Then we calculate Kendall’s rank correlation coefficient [46], if the ranking on the vertices mismatch.
@@ 6.1 Comparison between PageRank and qPageRank for a few simple graphs
(a) The Path graph with $60$ vertices.
@@
Figure 1: Comparison between the PageRank and qPageRank of vertices on path graph with $60$ vertices. (Color online)
@@ 6.2 Comparison between PageRank and qPageRank for a few directed graphs
Now, we consider the directed graphs, where most edges have only one direction. There are a few graphs for which PageRank and qPageRank act similarly, such as the directed trinary tree, GN, GNC, and GNR networks [50]. For many others PageRank and qPageRank behave differently, for instance, the scale-free network, Erdős-Rényi random graph, random_k_out graph, etc. We initiate our discussion with the graphs for which PageRank and qPageRank match.
We first compute our qPageRank for a directed graph depicted in Figure 3(a), which was also considered in [23] earlier. This graph does not belong to any particular class of graphs. We find that the ranking generated by DTOQW is different from the PageRank and the qPageRank developed in [23]. The values of DTOQW-based qPageRank for different vertices are available in Table 6. Also, Table 6 provides a comparison between different rankings. Different PageRank values for the vertices are represented by a multiple-bar diagram in Figure 3(b).
Vertices as labeled in Figure 3(a) qPageRank based on DTOQW Ranking based on DTOQW Ranking based on qPageRank developed in [23] Ranking based on PageRank [23]
1 0.0434 6 5 6
2 0.2895 1 4 5
3 0.0601 4 3 4
4 0.0272 7 7 7
5 0.2627 3 2 2
6 0.0473 5 6 3
7 0.2697 2 1 1
Table 6: PageRank and qPageRank values for a toy graph depicted in Figure 3(a).
(a) A directed graph which does not belong to any specific class of graphs, considered in [23].
Figure 3: We consider a directed graph with $7$ vertices and $11$ edges which are depicted in subfigure 3(a). For the vertices we plot the PageRank, qPageRank based on [23] and qPageRank based on DTOQW with bar diagrams in subfigure 3(b). (Color online.)
Now, we consider a directed trinary tree for our computation. Here, edges are directed outwards from a central vertex 00. The vertices are distributed into four layers surrounding the central vertex. The vertices in the same layer have equal PageRank and qPageRank, which we mention in Table 7.
@@
We plot the values of PageRank and qPageRank of the directed tree using bar diagrams. The graph and its corresponding bar diagrams are kept in Figure 4. The PageRank and qPageRank indicate the same ranking pattern on the directed trinary tree. As a result, the rank correlation is $1$. Also, the difference in PageRank of the vertices in the third and fourth layers is $0.000039$. But, the difference in qPageRank of the vertices in the third and fourth layers is .0091.0091.0091.0091. Therefore, the qPageRank is more sensitive for determining the vertices in the outer layer. Also, compare the bar diagrams in Figures 2(b) and 4(b), which are depicted for the undirected and directed trinary trees with an equal number of vertices, respectively. The qPageRank is more efficient than the classical PageRank for finding the vertices in the outermost layer in the case of a directed trinary tree.
@@
The Erdős-Rényi random graph [53] is a family of well-known random graphs. The edges are selected randomly to construct a random graph. We consider a specific instance of a directed Erdős-Rényi random graph with $80$ vertices and $283$ edges, which we plot in figure 9(a). In this case, the PageRank and qPageRank behave differently. The top five vertices with the highest PageRank are $5,79,78,45$, and $3$. Also, the top five vertices with the highest qPageRank are $26,67,77,68$, and $17$. There is no common vertex. The difference between PageRank and qPageRank for the Erdős-Rényi random graph is also talked about in other qPageRank proposals [23, 26].
@@ 6.3 Convergence test of qPageRank
For our numerical experiment to determine convergence of qPageRank procedure, we consider a random sample of the vertices from a graph, which includes the vertex with the maximum qPageRank. For different time instance $t=0,1,2,\dots T$, we calculate $q\Pi^{(t)}$ and plot corresponding values for distinct vertices in our sample. Note that, we consider accuracy of $q\Pi$ upto three decimal places. Therefore, $q\Pi^{(T)}$ represents the accurate value of $q\Pi$ upto three decimal places. For better accuracy the number of iterations may be increased. For different graphs we collect the plots in Figure 11.
@@ 6.4 Dependency of qPageRank on $\alpha$
There are experimental evidences showing that the PageRank exhibits a dependency on the damping parameter $\alpha$. In fact, the choice of $\alpha$ in the classical PageRank algorithm is also motivated by the need to ensure fast convergence of the power method. It ensures a reasonably large spectral gap in the Google matrix. The usual choice of $\alpha=0.85$ is expected to produce a satisfactory web surfing model, as well as a good convergence rate in computing the PageRank vector. The stability of the qPageRank is also a crucial factor to consider, as it randomly sets the damping parameter to a certain value. In fact, there is no a priori argument available to set the value of $\alpha$. The initial value of $0.85$ in the conventional PageRank protocol was selected to resemble the actions of a random walker. Given the arbitrary selection of the parameter $\alpha$, it is highly preferred that the algorithm’s output remain stable. Hence, it is interesting to investigate how the qPageRank is inspired by the choice of $\alpha$. For the other proposals, the dependency of qPageRank on $\alpha$ is investigated in [21, 23, 24, 26, 17]
@@
We calculate the trace distance $D(\rho_{average},\rho^{(T)})$ and fidelity $F(\rho_{average},\rho^{(T)})$ between the average state $\rho_{average}$ and the final state $\rho^{(T)}$ for different values of $\alpha$. In figure 12 and 13, we plot $D(\rho_{average},\rho^{(T)})$ and $F(\rho_{average},\rho^{(T)})$ with respect to $\alpha$ for different undirected and directed graphs, respectively, under our consideration. Below we summaries our observations from these figures.
Consider the undirected graphs first. We observe that if the trace distance $D(\rho_{average},\rho^{(T)})$ has a local maxima for some $\alpha=\alpha_{0}\in(0,1)$, then the fidelity $F(\rho_{average},\rho^{(T)})$ has a local minima at $\alpha_{0}$. For instance, consider the graph for $D(\rho_{average},\rho^{(T)})$ and $F(\rho_{average},\rho^{(T)})$, which are depicted in sub-figures 12(a) and 12(b), respectively, for the balanced tree $(3,4)$. In sub-figure 12(a) we observe a local maxima at $\alpha=0.85$. Also, in sub-figure 12(b) we obtain a local minima at $\alpha=0.85$. Similarly, these local maxima and minima can be observed at $\alpha=0.75$ in the case of Karate club network, which are depicted in figures 12(c) and 12(d). The Barabási and Albert network has local maxima and minima for $\alpha=0.95$, which is depicted in sub-figures 12(e) and 12(f). In the case of the Watts-Strogatz network there are no such local maxima or minima. The corresponding curves are drawn in sub-figure 12(g) and 12(h). Therefore, we may assume that $\alpha=0.85$ is a wise choice.